OpenStudy (johnw):

f(x) = arccos(x^2+1) What´s the derivative of the function?

OpenStudy (holsteremission):

Do you know the chain rule? Do you know the derivative of the inverse cosine function?

OpenStudy (johnw):

Well, the derivative of arccos is -1/√1-x^2, right?

OpenStudy (sshayer):

\[\frac{ d }{ dx }(\cos^{-1} u)=\frac{ -1 }{ \sqrt{1-u^2} }u \prime \] where u is a function of x

OpenStudy (dumbcow):

yes follow what the above post says. the substitution is that u = x^2 + 1 du = 2x

OpenStudy (danjs):

Using that substitution \[u = x^2 + 1\] \[\frac{ du }{ dx }=2x\] the chain rule is \[\frac{ df }{ dx }=\frac{ df }{ du }*\frac{ du }{ dx }\]

OpenStudy (danjs):

\[f(u)=\cos^{-1} (u)\] \[\frac{ df }{ dx }=\frac{ df }{ du }*\frac{ du }{ dx } = \frac{ -1 }{\sqrt{ 1-u^2} }*\frac{ du }{ dx }\]

OpenStudy (danjs):

just replace back to f(x) for u and du/dx

OpenStudy (danjs):

\[\frac{ df }{ dx }=\frac{ -1 }{ \sqrt{1-(x^2 + 1)^2} } * 2x\]

OpenStudy (danjs):

not sure if that root can be a real number though